OpenStudy (anonymous):

Decide which part of the quadratic formula tells you whether the quadratic equation can be solved by factoring. −b b2 − 4ac 2a Use the part of the quadratic formula that you chose above and find its value given the following quadratic equation: 4x2 + 6x + 2 = 0

OpenStudy (ranga):

(b^2 - 4ac) is called the discriminant. If discriminant is positive and a perfect square then the quadratic equation can be solved by factoring.

OpenStudy (anonymous):

okay thanks so is it (2x+1)(x+3). as the value

OpenStudy (anonymous):

@ranga

OpenStudy (ranga):

4x2 + 6x + 2 = 0 First let us find the discriminant: b^2 - 4ac 6^2 - (4)(4)(2) = 36 - 32 = 4 Discriminant is positive and a perfect square and therefore the quadratic equation can be solved by factoring. Here I don't think they are asking you to factor it. They are saying: "Use the part of the quadratic formula that you chose above and find its value" The part you chose for the previous question was: b^2 - 4ac and its value is 4. But if you need to factor it then: ...

OpenStudy (ranga):

4x^2 + 6x + 2 = 0 4x^2 + 4x + 2x + 2 = 0 4x(x + 1) + 2(x + 1) = 0 (4x + 2)(x + 1) = 0 we can further factor 2 out of the first factor: 2(2x + 1)(x + 1) = 0

OpenStudy (anonymous):

thank you for your help!

OpenStudy (ranga):

you are welcome.

OpenStudy (anonymous):

@ranga so what is the answer